# Addressing outliers in mixed-effects logistic regression: a more robust modeling approach

### v1

@@ Abstract
This study introduces an outlier-robust model for analyzing hierarchically structured bounded count data within a Bayesian framework, utilizing a logistic regression approach implemented in JAGS. Our model incorporates a $t$-distributed latent variable to address overdispersion and outliers, improving robustness compared to conventional models such as the beta-binomial, binomial-logit-normal, and standard binomial models. Notably, our approach models the median of the response variable, presenting a more convenient and interpretable measure of central tendency, which is available in closed form. For comparability between all models, we also make predictions based on the mean proportion; however, this involves an integration step for the $t$-distributed nuisance parameter. While limited literature specifically addresses outliers in mixed models for bounded count data, this research fills that gap. The practical utility of the model is demonstrated using a longitudinal medication adherence dataset, where patient behavior often results in abrupt changes and outliers within individual trajectories. A simulation study demonstrates the binomial-logit-$t$ model’s strong performance, with comparison statistics favoring it among the four evaluated models. An additional data contamination simulation confirms its robustness against outliers. Our robust approach maintains the integrity of the dataset, effectively handling outliers to provide more accurate and reliable parameter estimates.
@@ 1   Introduction
Our framework can also be interpreted as an observation-level random effects (OLRE) logistic regression designed for overdispersed bounded count data [12] within mixed-effects models, incorporating both cluster-specific random effects and OLREs. OLRE models introduce random effects at the level of individual observations, unlike conventional mixed-effects models that typically include random effects at the group or cluster level. Unlike traditional OLRE models that often use Gaussian distributions to handle overdispersion, our approach incorporates a $t$-distributed latent variable as the OLRE, providing robustness to outliers in addition to overdispersion. A key advantage of our model is that the link function of covariates is directly related to the marginal median unconditional on the OLREs, making interpretation more convenient. Additionally, the median is often considered a better measure of central tendency when the data are skewed, as it is less influenced by extreme values compared to the mean. However, for the mean response, one must integrate out the latent variable, adding complexity. Nonetheless, our method distinctively extends the capabilities of OLRE marginalization techniques by applying a $t$-distributed framework, thereby offering enhanced robustness to outliers.
@@ 2.1 Dataset and methodological overview
For a more comprehensive discussion of the data from Vrijens et al. [24], including specifics on the ‘adherence’ variable, readers are referred to Burger et al. [9], which provides details on the aspects of the dataset not extensively covered in this study.
@@ 2.2 Challenges in modeling adherence data
  • The Kumaraswamy distribution lacks robustness against outliers, necessitating adopting an approach that can effectively manage such outliers [8].
@@ 3   Binomial-logit-t distribution
where $t\left(\xi_{i}\left|\eta,\sigma^{2},\nu\right.\right)$ is the probability density function (PDF) of the $t$-distribution with mean $\eta$, scale parameter $\sigma^{2}$, and degrees of freedom $\nu$. Integrating out the nuisance parameter $\xi_{i}$ leads to a distribution with heavier tails. This property improves the model’s robustness, effectively handling outlier values that are otherwise influential in lighter-tailed distributions.
The median of $Y_{i}$ can be directly linked to the parameter $\eta$, and is given by
$\text{Median}\left(Y_{i}\right)=m_{i}\frac{1}{1+e^{-\eta}}.$ (2)
This specific form for the median arises because the logistic function is applied to the median of the latent variable $\xi_{i}$, directly determining the median of $Y_{i}$ by $\eta$. This expression is convenient as it avoids the need for numerical integration, simplifying the interpretation of regression parameters. The underlying logic is that for a monotonic transformation, the median of the transformed variable is the transformation of the median of the original variable. While medians are typically integers, this continuous approximation is used to simplify model interpretation and application.
@@ 4   Mixed-effects models for bounded counts
In the following subsections, we present the series of mixed-effects models for bounded counts, starting with the simplest and least robust models and progressively introducing more complex assumptions. Table 1 summarizes the key characteristics of each model, including their ability to accommodate overdispersion and outliers and their capacity to model the mean and median response proportions.
Table 1: Comparison of mixed-effects models for bounded counts.
@@
Mixed model Overdispersion Outliers Mean Median
Binomial No No No (closed-form) Same as mean
Beta-binomial Yes No No (closed-form) Not in closed-form
@@ 4.1 Binomial model
The mean and median of the binomial distribution are equivalent, both directly determined by the probability of success.
@@ 4.2 Beta-binomial model
$P\left(y_{ij}\left|\bm{\beta},\bm{u}_{i},\delta\right.\right)=\int_{0}^{1}\binom{m_{ij}}{y_{ij}}\pi_{ij}^{y_{ij}}(1-\pi_{ij})^{m_{ij}-y_{ij}}\text{Beta}\left(\pi_{ij}\left|\theta_{ij},\gamma_{ij}\right.\right)\mathrm{d}\pi_{ij},$ (8)
where $\text{Beta}\left(\pi_{ij}\left|\theta_{ij},\gamma_{ij}\right.\right)$ denotes the PDF of the beta distribution. The shape parameters $\theta_{ij}=p_{ij}\delta$ and $\gamma_{ij}=\left(1-p_{ij}\right)\delta$ are derived from the logistic regression of the probability of success $p_{ij}$, as shown in Equation (6).
@@
The median of the beta-binomial distribution is not available in closed form and requires numerical methods to estimate.
@@ 4.3 Binomial-logit-normal model
where $\eta_{ij}=\bm{x}_{ij}^{\prime}\bm{\beta}+\bm{z}_{ij}^{\prime}\bm{u}_{i}$ represents the mean of the normal distribution, and $\xi_{ij}$ are the nuisance parameters that are integrated out. The term $\mathcal{N}\left(\xi_{ij}\left|\eta_{ij},\sigma^{2}\right.\right)$ denotes the PDF of the normal distribution.
@@
The median of the binomial-logit-normal model is directly linked to the linear predictor without the need for integration. Specifically, the median value is given by
$\text{Median}\left[y_{ij}\left|\bm{u}_{i}=\bm{0}\right.\right]=m_{ij}\frac{1}{1+e^{-\bm{x}_{ij}^{\prime}\bm{\beta}}}.$ (11)
@@
As a special case, when $\sigma^{2}$ approaches zero, the model simplifies to the conventional binomial distribution. In this scenario, the expected value calculation, $\mathbb{E}\left[y_{ij}\left|\bm{u}_{i}=\bm{0}\right.\right]$, reduces to the expression for the median in Equation (11), similar to the expected value in standard logistic regression models.
@@ 4.4 Binomial-logit-t model
Similarly, the median value of $y_{ij}$ can be directly linked to the linear predictor without the need for integration as in Equation (11).
@@ 5   Model comparison and adequacy
This section discusses model evaluation using three principal metrics. We employ the WAIC to compare models, assessing both fit and complexity [26] with the R package loo [23]. The K-L divergence is used to evaluate model adequacy, identifying influential observations and outliers [25, 22, 17], as described in Appendix B. Additionally, we assess model residuals as a measure of goodness of fit based on the posterior predictive distribution (see Appendix C[13]. We calculated p-values for uniformity (Kolmogorov-Smirnov test), overdispersion (parametric bootstrap test), and outliers (binomial test) based on residuals obtained through posterior predictive checks, providing quantitative diagnostics for model adequacy.
@@ 6.1 Model implementation
We modeled the mean proportion under all four models and the median with the binomial-logit-normal and binomial-logit-$t$ models to compare model fits for both central tendency measures.
@@
We estimated the adherence proportions over time for each intervention group under each regression model. For each posterior sample, adherence proportions were calculated for a series of time points across the different intervention groups. Specifically, we computed the predicted probabilities of adherence for each combination of time (from months 0 to 10) and intervention group (intervention and non-intervention), setting the random effects to zero to represent the average patient and using the average age as a covariate. For the binomial-logit-normal and binomial-logit-$t$ models, the calculated mean probabilities are not conditional on the distributions’ latent (or nuisance) parameters $\xi_{ij}$, aligning with Equations (13) and (10), where expected values are computed independent of $\xi_{ij}$. Additionally, we computed the median probabilities for the binomial-logit-normal and binomial-logit-$t$ models directly from the linear predictors, as described in Equation (11), avoiding the need for integration.
We summarized the posterior results by calculating the median of the predicted mean probabilities for each time point and intervention group, serving as estimates for mean adherence proportions. For the binomial-logit-normal and binomial-logit-$t$ models, we also calculated the median of predicted median probabilities to estimate median adherence proportions. Additionally, 95% HPD intervals were computed to quantify the uncertainty of each estimate.
@@ 6.2 Regression fits and model comparison
In addition to evaluating mean adherence, we assessed median adherence proportions. This analysis examines the median adherence estimates between the binomial-logit-normal and binomial-logit-$t$ models. The median adherence proportions and their corresponding 95% HPD intervals for the intervention and non-intervention groups over the 10-month period are presented in Figure 8.
@@
  • The 95% HPD intervals of the median adherence proportions for the non-intervention group are slightly wider under the binomial-logit-normal model compared to the binomial-logit-$t$ model, reflecting greater uncertainty. In contrast, the adherence estimates for the intervention group exhibit similar levels of uncertainty between the two models.
@@
Figure 8: Median adherence proportion estimates and 95% highest posterior density (HPD) intervals under the binomial-logit-normal and binomial-logit-$t$ models over a 10-month period. The left panel shows non-intervention patients, while the right panel shows intervention patients. Gray dotted lines with diamond markers represent the binomial-logit-normal model, and dodgerblue solid lines with triangle markers represent the binomial-logit-$t$ model. The results show that the 95% HPD intervals for the non-intervention group are slightly wider for the binomial-logit-normal model. For the intervention group, the results are similar between the two models.
@@ 6.3 Model adequacy
Binomial-logit-normal 0.032 0.218 $<0.001$
Beta-binomial $<0.001$ 0.140 $<0.001$
Binomial $<0.001$ 0.151 $<0.001$
@@ 8   Discussion
A significant advantage of our approach is the availability of a closed-form solution for the median of the binomial-logit-$t$ distribution, which contrasts with the model proposed by Bayes et al. [5] that provides a closed-form expression for the mean. This simplification facilitates a more straightforward interpretation of the regression parameters, especially in the context of skewed data, where the median is often considered a more appropriate measure of central tendency than the mean. Initially, we attempted to implement Bayes et al.’s [5] model using a Bayesian framework and extend it to a mixed model setting. However, we encountered convergence issues, indicating the potential computational challenges of their approach. Despite these issues, our model shows adequate performance in terms of bias and coverage, demonstrating its effectiveness in analyzing overdispersed and heavy-tailed count data. For the mean proportion predictions, each integration per posterior iteration requires numerical integration to account for the latent variables. Our code runs efficiently by utilizing R’s built-in functions to specify the binomial PMF and $t$-distribution’s PDF, simplifying the integral calculation. However, integrating the latent variable for mean predictions points to an opportunity for future improvements and enhancements in our method.
The specified multivariate normal distribution for the cluster-specific random effects, $\bm{u}_{i}$, is not robust to outliers at the cluster level; however, it maintains robustness to outliers on a within-cluster basis. In order to improve robustness against outliers in cluster-specific effects, the model could be extended by adopting a multivariate $t$-distribution for the random effects. For greater flexibility in modeling, one could assume a skew-$t$ distribution for the latent variables or extend the approach to a more generalized family of skew-elliptical distributions, which would allow for the modeling of both asymmetry and heavier tails.
@@
The structural differences between our models, specifically between the binomial and beta-binomial models and the binomial-logit-normal and binomial-logit-$t$ models, influence their comparability when using the WAIC for model assessment. Unlike the binomial and beta-binomial models that directly link the probability of success to the covariates and predict the mean response, the binomial-logit-normal and binomial-logit-$t$ models link the median response to the covariates. This fundamental difference in the target of estimation adds complexity to the models’ structures, which could affect the WAIC’s ability to compare model fit and complexity fairly across different model types. This discrepancy is important to consider when interpreting the WAIC results, as it may lead to biased comparisons due to the different targets (mean vs. median).
In conclusion, the binomial-logit-$t$ model offers a robust solution for analyzing bounded count data with overdispersion and outliers. A key advantage of our model is the closed-form expression for the median, providing a straightforward interpretation of regression parameters. Although estimating the mean involves numerical integration, we successfully model both the mean and median proportions, making our approach versatile in mixed modeling frameworks. The strong simulation performance and demonstrated robustness in real-world applications confirm the practical effectiveness of our model. Future work should focus on optimizing computational methods and extending the model to accommodate a broader range of data structures.

### v2

@@ Abstract
This study introduces an outlier-robust model for analyzing hierarchically structured bounded count data within a Bayesian framework, utilizing a logistic regression approach implemented in JAGS. Our model incorporates a $t$-distributed latent variable to address overdispersion and outliers, improving robustness compared to conventional models such as the beta-binomial, binomial-logit-normal, and standard binomial models. Notably, our model targets a pseudo-median that differs from the true discrete median by less than one count; this closed-form quantity provides a robust and interpretable measure of central tendency. For comparability between all models, we additionally make predictions based on the mean proportion; however, this involves an integration step for the $t$-distributed nuisance parameter. While limited literature specifically addresses outliers in mixed models for bounded count data, this research fills that gap. The practical utility of the model is demonstrated using a longitudinal medication adherence dataset, where patient behavior often results in abrupt changes and outliers within individual trajectories. A simulation study demonstrates the binomial-logit-$t$ model’s strong performance, with comparison statistics favoring it among the four evaluated models. An additional data contamination simulation confirms its robustness against outliers. Our robust approach maintains the integrity of the dataset, effectively handling outliers to provide more accurate and reliable parameter estimates.
@@ 1  Introduction
Our framework can also be interpreted as an observation-level random effects (OLRE) logistic regression designed for overdispersed bounded count data [15] within mixed-effects models, incorporating both cluster-specific random effects and OLREs. OLRE models introduce random effects at the level of individual observations, unlike conventional mixed-effects models that typically include random effects at the group or cluster level. Unlike traditional OLRE models that often use Gaussian distributions to handle overdispersion, our approach incorporates a $t$-distributed latent variable as the OLRE, providing robustness to outliers in addition to overdispersion. A key advantage of our model is that the link function of covariates is directly related to the marginal pseudo-median (within one count of the true discrete median), unconditional on the OLREs, making interpretation more convenient. Because the pseudo-median is less influenced by extremes, it is often a better summary of central tendency in the presence of outliers. Additionally, the (pseudo-)median is often considered a better measure of central tendency for skewed data. However, for the mean response, one must integrate out the latent variable, adding complexity. Nonetheless, our method distinctively extends the capabilities of OLRE marginalization techniques by applying a $t$-distributed framework, thereby providing improved robustness to outliers.
@@ 2.1 Dataset and methodological overview
For a more comprehensive discussion of the data from Vrijens et al. [28], including specifics on the “adherence" variable, readers are referred to Burger et al. [9], which provides details on the aspects of the dataset not extensively covered in this study.
@@ 2.2 Challenges in modeling adherence data
  • The Kumaraswamy distribution lacks robustness against outliers, necessitating the adoption of an approach that can effectively manage such outliers [8].
@@ 3  Binomial-logit-t distribution
[…]
A convenient measure of location is obtained by passing the location parameter $\eta$ through the inverse logit and scaling by $m_{i}$:
$\widetilde{M}_{i}=m_{i}\text{logit}^{-1}\left(\eta\right).$ (3)
Because $\eta$ is the median of $\xi_{i}$ and the logistic function is strictly increasing, $\text{logit}^{-1}\left(\eta\right)$ is the median of the latent probability $p_{i}$. Hence, Equation (3) acts as a pseudo-median for the mixture-distributed count $Y_{i}$. It is not, in general, the exact discrete median, but the appendix shows:
$k_{i-}\leq\widetilde{M}_{i}<k_{i-}+1,$ (4)
where $k_{i-}$ is the smallest integer with $P\left(Y_{i}\leq k_{i-}\right)\geq 0.5$. We, therefore, refer to $\widetilde{M}_{i}$ as a “pseudo-median" of $Y_{i}$, because it preserves the latent median on the probability scale while forgoing the integer constraint on $Y_{i}$. We adopt $\widetilde{M}_{i}$ throughout for interpreting the fixed-effects location, since it remains both closed-form and numerically faithful.
More generally, any latent distribution whose median equals $\eta$ will yield the same pseudo-median property; the $t$ family is adopted here solely for its adjustable heavy tails.
@@ 4  Mixed-effects models for bounded counts
In the following subsections, we present a series of mixed-effects models for bounded counts, starting with the simplest and least robust models and progressively introducing more complex assumptions. Table 1 summarizes the key characteristics of each model, including their ability to accommodate overdispersion and outliers and their capacity to model the mean and pseudo-median response proportions.
Table 1: Comparison of mixed-effects models for bounded counts. “Integration required" refers to whether numerical integration is needed to obtain the mean or median (pseudo-median) on the original scale.
@@
Mixed model Overdispersion Outliers Mean Median (pseudo-median)
Binomial No No No (closed-form) No (same as mean)
Beta-binomial Yes No No (closed-form) Yes (no closed-form)
@@ 4.1 Binomial model
Because there is no mixture component, the (pseudo-)median of the binomial distribution equals its mean, both determined by $p_{ij}$.
@@ 4.2 Beta-binomial model
$P\left(y_{ij}\left|\bm{\beta},\bm{u}_{i},\delta\right.\right)=\int_{0}^{1}\binom{m_{ij}}{y_{ij}}p_{ij}^{y_{ij}}(1-p_{ij})^{m_{ij}-y_{ij}}\text{Beta}\left(p_{ij}\left|\theta_{ij},\gamma_{ij}\right.\right)\mathrm{d}p_{ij},$ (10)
[…]
@@
The pseudo-median of the beta-binomial distribution is not available in closed form and requires numerical methods to estimate.
@@ 4.3 Binomial-logit-normal model
[…]
@@
The pseudo-median of the binomial-logit-normal model is directly linked to the linear predictor without the need for integration. Specifically, the pseudo-median value is given by
$\widetilde{M}\left[y_{ij}\left|\bm{u}_{i}=\bm{0}\right.\right]=m_{ij}\frac{1}{1+e^{-\bm{x}_{ij}^{\prime}\bm{\beta}}}.$ (13)
@@
[…]
@@ 4.4 Binomial-logit-t model
Similarly, the pseudo-median value of $y_{ij}$ can be directly linked to the linear predictor without the need for integration as in Equation (13).
@@ 5  Model comparison and adequacy
We evaluate model performance using three complementary criteria.
First, we apply the WAIC [30], calculated using the loo package [27]. WAIC balances model fit and complexity via
$\text{WAIC}=\operatorname{lppd}-p_{\text{WAIC}},$ (16)
where $\operatorname{lppd}$ is the log pointwise predictive density and $p_{\text{WAIC}}$ is the effective number of parameters. Lower WAIC values indicate better expected out-of-sample performance.
Second, we approximate the K-L divergence between each model’s posterior predictive distribution and the empirical distribution using the leave-one-cluster-out importance-sampling estimator described in Appendix B [29, 26, 20]. Large K-L values flag influential observations and possible model misspecification.
Third, we assess goodness of fit using posterior predictive residual diagnostics from the DHARMa package [16]. We report Kolmogorov-Smirnov $p$-values for uniformity, parametric bootstrap $p$-values for overdispersion, and binomial $p$-values for outliers; $p$-values below 0.05 signal a lack of fit.
@@ 6.1 Model implementation
We modeled the mean proportion under all four models and the pseudo-median with the binomial-logit-normal and binomial-logit-$t$ models to compare model fits for both central tendency measures.
@@
We estimated the adherence proportions over time for each intervention group under each regression model. For each posterior sample, adherence proportions were calculated for a series of time points across the different intervention groups. Specifically, we computed the predicted probabilities of adherence for each combination of time (from months 0 to 10) and intervention group (intervention and non-intervention), setting the random effects to zero to represent the average patient and using the average age as a covariate. For the binomial-logit-normal and binomial-logit-$t$ models, the calculated mean probabilities are not conditional on the distributions’ latent (or nuisance) parameters $\xi_{ij}$, aligning with Equations (15) and (12), where expected values are computed independent of $\xi_{ij}$. Additionally, we computed the closed-form pseudo-median probabilities for the binomial-logit-normal and binomial-logit-$t$ models directly from the linear predictors (Equation (13)), avoiding the need for integration and thereby staying close to the discrete median.
We summarized the posterior results by calculating the median of the predicted mean probabilities for each time point and intervention group, serving as estimates for mean adherence proportions. For the binomial-logit-normal and binomial-logit-$t$ models, we also calculated the median of predicted median probabilities to estimate pseudo-median adherence proportions. Additionally, 95% HPD intervals were computed to quantify the uncertainty of each estimate.
@@
All models were run on a 4-core Intel i5-1145G7 laptop (16 GB RAM). Convergence times were modest: binomial $\approx$ 5 min, beta-binomial $\approx$ 120 min, binomial-logit-normal $\approx$ 8 min, and binomial-logit-$t$ $\approx$ 20 min. Thus, the added robustness of the $t$ model incurs only a moderate computational cost.
On a separate simulated dataset mirroring the adherence-study design, we also fitted the binomial-logit-$t$ model in Stan using its default No-U-Turn Sampler (NUTS) [10]. Posterior summaries were similar to the JAGS results, and total run-time was of the same order. We, therefore, retain JAGS as the primary engine while making the Stan code available in the project’s GitHub repository for readers who prefer Hamiltonian Monte Carlo sampling.
@@ 6.2 Regression fits and model comparison
In addition to evaluating mean adherence, we assessed pseudo-median adherence proportions. This analysis examines the pseudo-median adherence estimates between the binomial-logit-normal and binomial-logit-$t$ models. The pseudo-median adherence proportions and their corresponding 95% HPD intervals for the intervention and non-intervention groups over the 10-month period are presented in Figure 8.
@@
  • The 95% HPD intervals of the pseudo-median adherence proportions for the non-intervention group are slightly wider under the binomial-logit-normal model compared to the binomial-logit-$t$ model, reflecting greater uncertainty. In contrast, the adherence estimates for the intervention group exhibit similar levels of uncertainty between the two models.
@@
Figure 8: Pseudo-median adherence proportion estimates and 95% highest posterior density (HPD) intervals under the binomial-logit-normal and binomial-logit-$t$ models over a 10-month period. The left panel shows non-intervention patients, while the right panel shows intervention patients. Gray dotted lines with diamond markers represent the binomial-logit-normal model, and dodgerblue solid lines with triangle markers represent the binomial-logit-$t$ model. The results show that the 95% HPD intervals for the non-intervention group are slightly wider for the binomial-logit-normal model. For the intervention group, the results are similar between the two models.
@@ 6.3 Model adequacy
Binomial-logit-normal 0.032 0.218 <0.001<0.001
Beta-binomial <0.001<0.001 0.140 <0.001<0.001
Binomial <0.001<0.001 0.151 <0.001<0.001
@@ 8  Discussion
A significant advantage of our approach is the availability of a closed-form solution for the pseudo-median of the binomial-logit-$t$ distribution, which contrasts with the model proposed by Bayes et al. [5] that provides a closed-form expression for the mean. This simplification facilitates a more straightforward interpretation of the regression parameters, especially in the context of skewed data, where the (pseudo-)median is often considered a more appropriate measure of central tendency than the mean. Initially, we attempted to implement Bayes et al.’s [5] model using a Bayesian framework and extend it to a mixed model setting. However, we encountered convergence issues, indicating the potential computational challenges of their approach. Despite these issues, our model shows adequate performance in terms of bias and coverage, demonstrating its effectiveness in analyzing overdispersed and heavy-tailed count data. For the mean proportion predictions, each integration per posterior iteration requires numerical integration to account for the latent variables. Our code runs efficiently by utilizing R’s built-in functions to specify the binomial PMF and $t$-distribution’s PDF, simplifying the integral calculation.
A Gibbs implementation using the Pólya-gamma augmentation of Polson et al. [24] is also possible; it would render the logistic part conditionally Gaussian and could improve mixing, but requires either a custom JAGS module or a different engine such as Stan.
The specified multivariate normal distribution for the cluster-specific random effects, $\bm{u}_{i}$, is not robust to outliers at the cluster level; however, the binomial-logit-$t$ model maintains robustness to outliers on a within-cluster basis. In order to improve robustness against outliers in cluster-specific effects, the model could be extended by adopting a multivariate $t$-distribution for the random effects.
[…]
@@
The structural differences between our models, specifically between the binomial and beta-binomial models and the binomial-logit-normal and binomial-logit-$t$ models, influence their comparability when using the WAIC for model assessment. Unlike the binomial and beta-binomial models that directly link the probability of success to the covariates and predict the mean response, the binomial-logit-normal and binomial-logit-$t$ models essentially link the pseudo-median of $Y$ (on the count scale) to the covariates, whereas the binomial/beta-binomial link the mean of $Y$ directly to the covariates. This fundamental difference in the target of estimation adds complexity to the models’ structures, which could affect the WAIC’s ability to compare model fit and complexity fairly across different model types. This discrepancy is important to consider when interpreting the WAIC results, as it may lead to biased comparisons due to the different targets (mean vs. pseudo-median).
In conclusion, the binomial-logit-$t$ model presents a robust solution for analyzing bounded count data with overdispersion and outliers. A key advantage of our model is the closed-form expression for the pseudo-median, providing a straightforward interpretation of regression parameters. Although estimating the mean involves numerical integration, we successfully model both the mean and pseudo-median proportions, making our approach versatile in mixed modeling frameworks. The strong simulation performance and demonstrated robustness in real-world applications confirm the practical effectiveness of our model. Future work should focus on optimizing computational methods and extending the model to accommodate a broader range of data structures.